Determine the length of a side of the rectangle whose perimeter is 52 inches and whose width is 1/2 its length.
Answer by checkley77(12844)

W=L/2
P=2L+2W
52=2L+2(L/2)
52=2L+L
3L=52
L=52/3
L=17.333 ANS.
W=17.333/2=8.667 ANS.
pROOF:
52=2*17.333+2*8.667
52=34.667+17.333
52=52